Sales Mix and Break-Even Sales

Data related to the expected sales of laptops and tablets for Tech Products Inc. for the current year, which is typical of recent years, are as follows:

Products Unit Selling Price Unit Variable Cost Sales Mix
Laptops $1,600 $800 40%
Tablets 900 450 60%

The estimated fixed costs for the current year are $3,894,000.

Required:

1. Determine the estimated units of sales of the overall (total) product, E, necessary to reach the break-even point for the current year.
______ units

2. Based on the break-even sales (units) in part (1), determine the unit sales of both laptops and tablets for the current year.

Laptops ____units
Tablets _____units

3. Assume that the sales mix was 60% laptops and 40% tablets. Compare the breakeven point with that in part (1). Why is it so different?
_____units
Answer #1
1
Weighted average unit contribution margin 590 =(1600-800)*40%+(900-450)*60%
Break-even point for the current year 6600 units =3894000/590
2
Laptops 2640 units =6600*40%
Tablets 3960 units =6600*60%
3
Weighted average unit contribution margin 660 =(1600-800)*60%+(900-450)*40%
Break-even point for the current year 5900 units =3894000/660
The break-even point is lower in this scenario than in part (1) because the sales mix is weighted more heavily toward the product with the higher contribution margin per unit of product.